Assume that Jose is indifferent between investing in a corporate bond that pays 10 percent interest and a stock with no growth potential that pays an 10 percent dividend yield. Assume that the tax rate on dividends is 15 percent. What is Jose's marginal tax rate? (Do not round intermediate computations.)

7.40 percent

20.40 percent

15.00 percent

None of the choices are correct.

30.20 percent
Answer #1

Answer:

Taxable dividend yield = 10%

Tax rate on dividend yield= 15%

Interest rate = 10%

Let tax rate on interest = X

Marginal tax rate

0.10*(1-X)=0.10*(1-0.15)

0.10-0.10X=0.10*0.85

0.10-0.10X= 0.085

0.10 - 0.085 = 0.10X

0.015 = 0.10X

X = 0.015/0.10

X= 0.15

X= 15%

Marginal tax rate= 15.00% (15.00 percent)